# Ratchet current and scaling properties in a nontwist mapping

### v1

@@ III Critical exponents and scaling law
The behavior of $I_{\mathrm{rms}}$ is shown in Fig. 8 for different perturbation values $\varepsilon$. The curves exhibit an accelerated growth regime for small $n$ whereas for large $n$, they achieve a saturation limit, characterized by a constant plateau due to the limited phase space area. The transition from accelerated growth to the constant plateau is given by a characteristic crossover number $n_{x}$. The behavior discussed above can be characterized by four critical exponents, namely, $\alpha$, $\beta$, $\gamma$, and $z$. To obtain them, we assume the following scaling hypotheses [89, 90, 91, 92, 75, 93, 94]
@@
Through numerical simulations, we obtain all these exponents. The acceleration exponent $\beta$ is obtained by plotting $I_{\mathrm{rms}}$ versus $n$ (Fig. 8) and performing a power law fitting $f(n)=Bn^{A}$ over the regime of growth. We find that the exponent $\beta$ corresponds to $A$, and equals $\beta=0.475\pm 0.008$  [^2] We performed the power law fitting for several values of $\varepsilon$ (those indicated in Fig. 9) and this value corresponds to $\expectationvalue{\beta}\pm\sigma_{\beta}$, where $\expectationvalue{\cdot}$ is the mean and $\sigma_{\beta}$ is the standard deviation., whereas $B=\varepsilon^{\alpha\beta}$. From different values of $\varepsilon$ we obtain different values of $B$ and from a power law fitting of $B$ versus $\varepsilon$ [Fig. 9(a)], we obtain $\alpha\beta=1.009\pm 0.009$, such that the transport exponent is $\alpha=2.12\pm 0.05$.
@@ IV Conclusions
We have explored the transport and diffusion of particles in the chaotic component in phase space in a newly reported nontwist Hamiltonian mapping. Firstly, our analysis focused on the properties of the survival probability considering two exits symmetrically placed on phase space with respect to the $I=0$ line. We have demonstrated that the survival probability follows an exponential decay when the chaotic sea dominates the survival region. For survival region defined with $I_{\mathrm{esc}}>I_{fi}$, where $I_{fi}$ is the center of the elliptical fixed point of the first period-1 island, the survival probability follows a stretched exponential with a power law tail for long times. Furthermore, we have shown that the decay rate scales with $I_{\mathrm{esc}}$ as a power law for $I_{\mathrm{esc}}<I_{fi}$, with exponent $\zeta=-1.865\pm 0.007$. By rescaling the horizontal axis of the plot $P(n)\times n$ by $n\rightarrow nI_{\mathrm{esc}}^{\zeta}$, we have found that the survival probability curves overlap into a single, and hence universal, plot for small survival regions, indicating that the survival probability maintains its behavior regardless of the size of the survival region.
@@
Secondly, we investigated whether the chaotic transport of particles has a preferential direction, i.e., whether the system exhibits unbiased transport. We have shown that considering particles distributed uniformly over the whole available phase space or particles randomly distributed over a small region around $I=0$, both ensembles initially with $\langle I\rangle\approx 0$, the tendency is for the particles to escape through the bottom exit, thus exhibiting the so-called ratchet effect. The mechanics responsible for this effect is an unbalanced stickiness due to asymmetry of the chaotic component in phase space with respect to the line $I=0$, i.e., the invariant spanning curves that bound the phase space are not symmetrical. This asymmetry generates a nonzero net current of particles due to different trapping times in the upper ($I>0$) and lower ($I<0$) regions of phase space. The cumulative distribution of recurrence times and the difference in the moments of the distribution of recurrence times support our claim for the existence of an unbalanced stickiness in phase space.
@@
Lastly, we presented a phenomenological description of diffusion in the chaotic component of phase space. We have chosen as our observable the square root of the averaged square action, $I_{\mathrm{rms}}$, and upon assuming three scaling hypotheses, we have found that the behavior of $I_{\mathrm{rms}}$ is characterized by four critical exponents. We have also derived an analytical scaling law relating these four exponents and from extensive numerical simulations, we have obtained all of them and showed that they remarkably agree with the scaling law. Our scaling hypotheses were supported by the collapse of the $I_{\mathrm{rms}}$ curves onto a single, and hence universal, curve.

### v2

@@ IV Critical exponents and scaling law
The behavior of $I_{\mathrm{rms}}$ is shown in Fig. 8 for different perturbation values $\varepsilon$. The curves exhibit an accelerated growth regime for small $n$ whereas for large $n$, they achieve a saturation limit, characterized by a constant plateau due to the bounded phase space area. The transition from accelerated growth to the constant plateau is given by a characteristic crossover number $n_{x}$. The behavior discussed above can be characterized by four critical exponents, namely, $\alpha$, $\beta$, $\gamma$, and $z$. To obtain them, we assume the following scaling hypotheses [96, 97, 98, 99, 81, 100, 101]
@@
Through numerical simulations, we obtain all these exponents. The acceleration exponent $\beta$ is obtained by plotting $I_{\mathrm{rms}}$ versus $n$ (Fig. 8) and performing a power law fitting $f(n)=Bn^{A}$ over the regime of growth. The exponent $\beta$ brings important information about diffusion. If $\beta=0.5$, normal diffusion dominates the dynamics. Exponents larger (smaller) than 0.5 means super (sub) diffusion. We find that the exponent $\beta$ corresponds to $A$, and equals $\beta=0.475\pm 0.008$  [^2] We performed the power law fitting for several values of $\varepsilon$ (those indicated in Fig. 9) and this value corresponds to $\expectationvalue{\beta}\pm\sigma_{\beta}$, where $\expectationvalue{\cdot}$ is the mean and $\sigma_{\beta}$ is the standard deviation., slightly less than 0.5, meaning the diffusion process is sub diffusive. From the same power law fitting, we find $B=\varepsilon^{\alpha\beta}$. From different values of $\varepsilon$ we obtain different values of $B$ and from a power law fitting of $B$ versus $\varepsilon$ [Fig. 9(a)], we obtain $\alpha\beta=1.009\pm 0.009$, such that the transport exponent is $\alpha=2.12\pm 0.05$.
@@ V Conclusions
The $\mathbf{E}\times\mathbf{B}$ drift motion is observed in many plasma devices and plays a major role in the anomalous transport of particles. In magnetic confinement devices, $\mathbf{E}\times\mathbf{B}$ shear is responsible for reducing turbulence levels, improving the confinement. Thus, understanding this drift motion is crucial for improving plasma confinement and optimizing various applications. In this paper, we have explored the transport and diffusion of particles in the chaotic component in phase space in a newly reported nontwist Hamiltonian mapping, that describes the advected particle motion under the $\mathbf{E}\times\mathbf{B}$ drift. Firstly, our analysis focused on the properties of the survival probability considering two exits symmetrically placed on phase space with respect to the $I=0$ line. We have demonstrated that the survival probability follows an exponential decay when the chaotic sea dominates the survival region. For survival region defined with $I_{\mathrm{esc}}>I_{fi}$, where $I_{fi}$ is the center of the elliptical fixed point of the first period-1 island, the survival probability follows a stretched exponential with a power law tail for long times. Furthermore, we have shown that the decay rate scales with $I_{\mathrm{esc}}$ as a power law for $I_{\mathrm{esc}}<I_{fi}$, with exponent $\zeta=-1.865\pm 0.007$. By rescaling the horizontal axis of the plot $P(n)\times n$ by $n\rightarrow nI_{\mathrm{esc}}^{\zeta}$, we have found that the survival probability curves overlap into a single, and hence universal, plot for small survival regions, indicating that the survival probability maintains its behavior regardless of the size of the survival region.
@@
Secondly, we investigated whether the chaotic transport of particles has a preferential direction, i.e., whether the system exhibits unbiased transport. We have shown that considering particles distributed uniformly over the whole available phase space or particles randomly distributed over a small region around $I=0$, both ensembles initially with $\langle I\rangle\approx 0$, the tendency is for the particles to escape through the bottom exit, thus exhibiting the so-called ratchet effect. The mechanics responsible for this effect is an unbalanced stickiness due to asymmetry of the chaotic component in phase space with respect to the line $I=0$, i.e., the invariant spanning curves that bound the phase space are not symmetrical. This asymmetry generates a nonzero net current of particles due to different trapping times in the upper ($I>0$) and lower ($I<0$) regions of phase space. The cumulative distribution of recurrence times and the difference in the moments of the distribution of recurrence times support our claim for the existence of an unbalanced stickiness in phase space. To the best of our knowledge, the existence of ratchet current in the $\mathbf{E}\times\mathbf{B}$ drift motion has not been reported in the literature. This fact has complex implications for the impurity transport, heating, and instabilities in the plasma, and future studies are necessary to better understand them.
@@
Lastly, we presented a phenomenological description of diffusion in the chaotic component of phase space. We have chosen as our observable the square root of the averaged square action, $I_{\mathrm{rms}}$, and upon assuming three scaling hypotheses, we have found that the behavior of $I_{\mathrm{rms}}$ is characterized by four critical exponents. One of the exponents, $\beta$, characterizes the diffusion process. We have obtained $\beta=0.475\pm 0.008$, which is slightly below 0.5, indicating that the diffusion process is subdiffusive. The subdiffusive nature of the system is related to the trappings that generate the stickiness effect a chaotic orbit experiences throughout its evolution [103]. We have also derived an analytical scaling law relating these four exponents and from extensive numerical simulations, we have obtained all of them and showed that they remarkably agree with the scaling law. Our scaling hypotheses were supported by the collapse of the $I_{\mathrm{rms}}$ curves onto a single, and hence universal, curve. Additionally, even though the obtained scaling law [Eq. (21)] is similar to the Family-Vicsek scaling discussed in Ref. [104], it is not directly connected. Nonetheless, this similarity motivates us to explore whether there is a correlation length that reaches the size of the bounded phase space area in the saturation regime. We intend to investigate this in a forthcoming paper.